OpenStudy (anonymous):

Can someone help with a few questions? Write y = –2/5x – 3 in standard form using integers. 2x + 5y = –15 2x + y = –3 2x + 5y = –3 x + y = –15 Tell whether the lines for the pair of equations are parallel, perpendicular, or neither. y = –2/3x + 1 2x – 3y = –3 parallel perpendicular neither Write the equation of a line that is perpendicular to the given line and that passes through the given point. y + 2 = 1/3(x – 5); (–4, 3) y = –3x + 15 y = –3x – 9 y = –4x + 3 y = –1/3x – 11 Thanks any help is appreciated :)

OpenStudy (anonymous):

is there something that you dont understand that can help you with

OpenStudy (anonymous):

All of it :\

OpenStudy (anonymous):

ok well then lets get started

OpenStudy (anonymous):

Ok so we can start with the first one? Write y = –2/5x – 3 in standard form using integers.

OpenStudy (anonymous):

yah just a sec

OpenStudy (anonymous):

sorry i forgot about question but i will get right on it

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

is this multple question or free answers

OpenStudy (anonymous):

multiple choice

OpenStudy (anonymous):

2x + 5y = –15 2x + y = –3 2x + 5y = –3 x + y = –15

OpenStudy (anonymous):

well i cantgive you the straight up answer but whatch this vido it might hel ou understand it just a bit better

OpenStudy (anonymous):

sorry my keyboard is so messed up

OpenStudy (anonymous):

what video?

OpenStudy (anonymous):

http://www.youtube.com/watch?v=47qv4w0A03s